OpenStudy (anonymous):

What is the slope of the line 6y + 2x = 18 in simplest form?

OpenStudy (princeharryyy):

-2/6 = -1/3

OpenStudy (zale101):

@Mathisnotfun! show your attempt on this problem so i can know were you went wrong.

OpenStudy (zale101):

Did you try to set up the equation 6y + 2x = 18 into the slope intercept form y=mx+b?

OpenStudy (anonymous):

I just have problems with all of these type of math problems. I simply don't get any of it.

OpenStudy (zale101):

The question wants you to find the slope of this equation 6y + 2x = 18. Once you set it up to y=mx+b where m is the slope. To do that, you try to solve the equation for y. \(\large 6y + 2x = 18\) subtract 2x to both sides to isolate 6y \(\large 6y + 2x-2x = 18-2x\) \(\large 6y = 18-2x\) reorder the terms to get 6y=-2x+18

OpenStudy (zale101):

\(\large 6y = -2x+18\) Bow, you try to reduce y by dividing everything by 6 \(\large\frac{ 6y}{6} = \frac {-2x}{6}+\frac{18}{6}\) \(\large y = \frac {-2x}{6}+3\)

OpenStudy (zale101):

Now*

OpenStudy (zale101):

clear so far?

OpenStudy (zale101):

\(\large y=-\frac{2}{6}x+3\) is the slope intercept form of the equation 6y + 2x = 18.

OpenStudy (zale101):

I feel like i'm being ignored :(